if a number is subtracted from twice its reciprocal, The result is -23/5.What is the number.
Answer by nerdybill(7384)

if a number is subtracted from twice its reciprocal, The result is -23/5.What is the number.
.
Let x = the number
then
1/x = the reciprocal
.
x - 2(1/x) = -23/5
x - 2/x = -23/5
multiplying both sides by 5x:
5x^2 - 10 = -23x
5x^2 + 23x - 10 = 0
(5x-2)(x+5) = 0
x = {2/5, -5}
answer: -5